calculate the mass of H2 needed to react with 5.6 g of N2


1
Expert's answer
2021-06-14T13:35:28-0400

3H2 + N2 "\to" 2NH3

1 mole of N2 react with 3 moles of H2

5.6g of N2 means moles of N2 = 5.6/28.013 = 0.199 moles

So moles of H2 needed = 3×0.199 = 0.597mol

Mass of H2 needed = 0.597 × 1.00784 = 0.601g